OpenStudy (anonymous):

(1/x+1) + (1/(1+(1/x)) = ?

OpenStudy (aum):

\[ \frac{1}{x+1} + \frac{1}{1 + \frac 1x} \\ \text{Multiply top and bottom of second fraction by x:} \\ \frac{1}{x+1} + \frac{x}{x + 1} = \frac{1+x}{x+1} = 1\\ \]

OpenStudy (anonymous):

thanks!

OpenStudy (aum):

yw.